OpenStudy (zeesbrat3):

The equation of a circle is (x - 4)2 + (y + 1)2 = 9. Which graph best represents the circle?

OpenStudy (anonymous):

A circle of Radius 3, with the Center at M(4/-1)

OpenStudy (zeesbrat3):

thanks so much. Do you think you can help me with another?

OpenStudy (anonymous):

(-: Lets see, what's the question?

OpenStudy (zeesbrat3):

Look at the graph. Which equation best represents the circle in the graph?

OpenStudy (anonymous):

Well, judging by my sight, M is M(-1/2)

OpenStudy (anonymous):

The Radius is 5

OpenStudy (zeesbrat3):

so its either c or d

OpenStudy (zeesbrat3):

no its d

OpenStudy (anonymous):

so it should be \[ f(x,y)= (x+1)^2+(y-2)^2=25\]

OpenStudy (zeesbrat3):

got them switched... its a

OpenStudy (anonymous):

ok, well this is the equation of the circle above.

OpenStudy (zeesbrat3):

thanks. How do you find the equations?

OpenStudy (anonymous):

the general solution of a circle looks like this \[ \large f(x,y)=(x-u)^2+(y-v)^2=r^2\]

OpenStudy (anonymous):

where \( u, v\) are the coordinates of the center.

OpenStudy (anonymous):

r = the radius, but it's always squared, so when you have 25, your radius is 5, when you have 9, your actual radius is 3, and so on and so forth.

OpenStudy (zeesbrat3):

Okay, thanks so much!

OpenStudy (anonymous):

very welcome